OpenStudy (anonymous):

Solve the system of equations x + y = 5 x - y = -9

OpenStudy (anonymous):

x=5-y 5-y-y=-9 -2y=-14 y=7

OpenStudy (anonymous):

add them 2x = -4 x=-2 subtract them 2y = 14 y=7 x=-2 , y=7

OpenStudy (anonymous):

Thanks